OpenStudy (anonymous):

How do you work this problem on a scientific calculator? a=5,000e (.082)(12)

OpenStudy (llort):

is e a variable?

OpenStudy (anonymous):

The problem is a=Pe exponent rt

OpenStudy (anonymous):

We are just not sure what buttons to push

OpenStudy (anonymous):

Do you mean: \[5000 \times e ^{0.82 \times 12}\]

OpenStudy (llort):

13373.9 there should be an fancy "e" button on your calculator, it could be a second function (possibly second to sqrt)

OpenStudy (anonymous):

Yes Mooby

OpenStudy (anonymous):

What order do you input the problem?

OpenStudy (anonymous):

5000 x e^(0.82 x 12) should do it.

OpenStudy (llort):

what make and model of calculator do you have?

OpenStudy (amistre64):

first you get a scientist, next ...

OpenStudy (anonymous):

Yeah my answer's like the TI-83 version.

OpenStudy (anonymous):

That is funny:) I will look at the calculator

OpenStudy (anonymous):

TI 30xa

OpenStudy (amistre64):

there should be an "e" button hidden on it someplace

OpenStudy (anonymous):

Yeah either e or y^x.

OpenStudy (anonymous):

When we push the ex button, it gives us an error

OpenStudy (amistre64):

e^x is asking for the exponent to it, fill in m*n depending on what m and n are

OpenStudy (amistre64):

i spose r*t in this case

OpenStudy (anonymous):

I just looked up the calculator on Google Images. Try using this button: \[y ^{x}\]

OpenStudy (llort):

not very many functions on your calculator... I believe it is the \[y^x\] button

OpenStudy (anonymous):

ok, so do we put 5000 x e?

OpenStudy (amistre64):

google does it up for you as well. 5,000*e^((.082)*(12))

OpenStudy (anonymous):

Oh, it does have an \[e ^{x}\]. Hrm.

OpenStudy (anonymous):

I pushed 5000x e and it says error

OpenStudy (anonymous):

Try doing it like this: 1. .82x12 2. e^x 3. x5000

OpenStudy (anonymous):

so, put .82x12xex5000

OpenStudy (llort):

did you try 2nd fxn LN?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I am just not sure what to do in between the steps. Like to do I mulitply them?

OpenStudy (anonymous):

.82x12 times e times 5000?

OpenStudy (llort):

e is a function, you need not multiply

OpenStudy (anonymous):

Yeah, I have TI-30X just realized there is no e key

OpenStudy (anonymous):

so should i put .82x12 then hit e and then multiply by 5000?

OpenStudy (anonymous):

13375.68 is the answer, but we just cannot get it on the cal

OpenStudy (amistre64):

what brand name you got?

OpenStudy (llort):

he has a TI 30xa

OpenStudy (anonymous):

Texas instruments

OpenStudy (anonymous):

http://mycalcdb.free.fr/galerie/TI/TI-30Xa.1.jpg

OpenStudy (anonymous):

this one?

OpenStudy (anonymous):

YES!

OpenStudy (llort):

seriously? there is a website devoted to pictures of calculators? **logs off of FantasyGirls** I know how I am going to spend my nights now.

OpenStudy (anonymous):

Ha ha ha

OpenStudy (amistre64):

2nd, ln, put in the r*t, hit " ) ", =

OpenStudy (amistre64):

or exactly: 2nd ln .082 x 12 ) =

OpenStudy (anonymous):

I come up with 0.984

OpenStudy (amistre64):

might wanna include the principal too tho :)

OpenStudy (anonymous):

where would we enter the 5000

OpenStudy (llort):

you can just multiply your answer... i think...

OpenStudy (amistre64):

afterwards just hit x 5000

OpenStudy (anonymous):

i have 49200 which is still not right

OpenStudy (amistre64):

in general: \begin{array}c 1.&5000\\ 2.&2^{nd}\\ 3.&ln\\ 4.&.082\\ 5.&x\\ 6.&12\\ 7.&)\\ 8.&=\\ \end{array}

OpenStudy (anonymous):

on step 3 I get an error

OpenStudy (amistre64):

maybe hit "x" between steps 1 and 2

OpenStudy (amistre64):

once you hit "ln", it should start itself off with a " ( ". if not, you might wanna include it

OpenStudy (anonymous):

still comes out to 0.984

OpenStudy (anonymous):

what about doing it on a TI-84 plus?

OpenStudy (cruffo):

probably need to enter the exponent first 0.82 times 12 2nd ln times 5000

OpenStudy (anonymous):

I got an error

OpenStudy (amistre64):

change the batteries :)

OpenStudy (anonymous):

Just did a few days ago:)

OpenStudy (cruffo):

For the TI-30a: probably need to enter the exponent first 0.082 times 12 2nd LN times 5000 equal For the TI-84 plus 5000 2nd LN 0.082 times 12 ) enter

OpenStudy (anonymous):

the other is solar

OpenStudy (anonymous):

YEA!!

OpenStudy (anonymous):

Cruffo got it!!

OpenStudy (anonymous):

.082*12 press"=" "2nd" "Ln" press"=" *5000 press "="

OpenStudy (anonymous):

it works on the graph

OpenStudy (anonymous):

THANK YOU GUYS!! GOD bless

OpenStudy (anonymous):

On a Hewlett Packard stack oriented calculator, for example the HP12C: 12 press ENTER . 082 press X (multiply) press "e^x" the result of e^(12 times .082) is left in the display. 5000 press X (multiply) Final result is in the the display, 13375.68

OpenStudy (anonymous):

Sorry. "the the display" should have been typed, "the display"

OpenStudy (cruffo):

RPN.... hummmmm